# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ 15 $\ln\left(\frac{1}{\alpha}\right)$ Dominated by the term 3
Table 2: terms from (B.23) and required conditions on $\left\lVert P^{\star}\mu\right\rVert_{2}$.
@@
$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k^{\star}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (B.24)
Once this requirement is satisfied, by Lemma B.17 and the inequality (B.23), we know that $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}P^{\star}X_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|$ will be less than $c_{1}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{0}$ and greater than $c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{1}$ for some $0<c_{1}<c_{2}<1$, and both with high probability. $c_{1}$ and $c_{2}$ are universal and their values can be determined following proper choices of $\epsilon,\alpha,N,k^{\star}$. From the argument at the end of the proof of Theorem 3.6, we know (B.24) leads to (3.4). The proof is completed. ∎
B.7 Proof of Corollary 3.12
@@
The techniques are parallel to the proof of the lower bounds above. By Lemma 3.15, for $k=k_{p,1}^{\star},c=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$, there exists a $\theta\in K$ such that $\left\lVert\theta\right\rVert_{p}=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{1}{(k_{p,1}^{\star})^{1/4}\sqrt{N}}\sigma$. Again, we impose a uniform prior on the set $\left\{\mu\left\lvert\right.\mu\in K,\mu_{i}=\kappa\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\in I\right\}$ for some positive constant $\kappa$. By the same arguments of Appendix B.2, if $\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$, or equivalently $\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}$ is sufficiently small, then there is no test with uniformly small type-I and type-II errors. We have
@@
and $\theta\in K$ from Lemma 3.15 satisfies $\left\lVert\theta\right\rVert_{p}=\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(Nk_{p,2}^{\star})^{1/4}}\sigma$. Consequently, it is not difficult to verify that $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\epsilon$. By Lemma B.1, we establish the second lower bound that $\rho_{\text{critical}}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$.
Suppose now that $\theta$ is the vector from Lemma 3.15 when $k=k_{p,3}^{\star}$. Subsequently, we know $\left\lVert\theta\right\rVert_{2}^{2}\leq\left\lVert\theta\right\rVert_{p}^{p}\cdot\left\lVert\theta\right\rVert_{\infty}^{2-p}\leq\epsilon^{2}\sigma^{2}$. Following the same reasoning in Appendix B.4, we know there always exists a strategy for $\mathcal{C}$ to discourage the testing. Therefore, we establish the third lower bound. The proof is completed. ∎
B.9 Proof of Theorem 3.18
@@
Lemma B.18 (Truncation of a $p$-convex orthosymmetric set).
@@
Lemma B.18 is straightforward from the definition. For any $\theta_{J,1},\theta_{J,2}\in K_{J}$, suppose $\theta_{J,1}=\beta_{1,J},\theta_{J,2}=\beta_{2,J},\beta_{1},\beta_{2}\in K$, then we know for any $\lambda\in[0,1]$, we have $\lambda\theta_{J,1}^{p}+(1-\lambda)\theta_{J,2}^{p}=\lambda\beta_{1,J}^{p}+(1-\lambda)\beta_{2,J}^{p}=(\lambda\beta_{1}^{p}+(1-\lambda)\beta_{2}^{p})_{J}\overset{\text{(i)}}{=}\widetilde{\beta}_{J}^{p}\in K_{J}^{p}$, where $\widetilde{\beta}\in K^{p}$ and (i) is from that $K$ is p-convex. Consequently, $K_{J}$ is p-convex.
@@
Suppose first $\epsilon\leq\frac{1}{\sqrt{N}}$. In this case, we have $\frac{j^{1/p-1/4}}{\sqrt{N}}\sigma\geq\frac{\sqrt{\epsilon}j^{1/p-1/4}}{N^{1/4}}\sigma$ for any $0\leq j\leq d$. By the monotonicity of Kolmogorov widths and the definitions of $k_{p,1}^{\star}$ and $k_{p,2}^{\star}$, we know $k_{p,1}^{\star}\leq k_{p,2}^{\star}$. Later calculations show that in this case we also have $k_{p,2}^{\star}\leq k_{p,3}^{\star}$. This implies $k_{p,1}^{\star}\leq k_{p,2}^{\star}\leq k_{p,3}^{\star}$. For $k_{p,1}^{\star}$ and its corresponding $P_{p,1}^{\star}$, without loss of generality we assume $P_{p,1}^{\star}\theta=(\theta_{1},\dots,\theta_{k_{p,1}^{\star}},0,\dots,0)$. Regard $P_{p,1}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,1}^{\star}}$, by the Hölder’s inequality A.17, we know $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ for any $\theta\in K$. Further define $k_{p,1,1}^{\star}:=\max\limits\left\{j\left\lvert\right.0\leq j\leq k_{p,1}^{\star},D_{j-1}(K)>\frac{j^{1/4}}{\sqrt{N}}\sigma\right\}$. Since $\epsilon\leq\frac{1}{\sqrt{N}}$, by the proof of Corollary 3.12 and Lemma B.18, we know that if $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, then there exists a desired test. However, this is already guaranteed since $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ and $k_{p,1}^{\star}\geq k_{p,1,1}^{\star}$.
@@
B.10 Verification of the Minimality in Section 4.2
@@
Appendix C Theoretical and Polynomial Time Algorithms
C.1 Theoretical Algorithm
According to Theorem 3.6 and Section B.5, we first attempt to search for a consistent set $T$ with respect to $\phi_{e}$. $\phi_{e}(\mathbf{X}_{T})$ then determines whether we are to accept or reject $H_{0}$.
1
2import $\mathbf{X},N,K,\sigma$
3
4Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma,\epsilon$ as 2.3 and 2.4
5if $k_{1}^{\star}\leq k_{2}^{\star}$ then
6 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
7
8else
9 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
10
[…]
13
14 consistent_flag == True
16    subset_test == True
17 else
18    subset_test == False
[…]
20
22       subsubset_test == True
23    else
24       subsubset_test == False
25    if subset_test XOR subsubset_test then
26       consistent_flag == False
27       break
28
29 if consistent_flag then
30    if subset_test then
31         Reject $H_{0}$
32       return
33    else
34         Accept $H_{0}$
35       return
36
37
38Accept $H_{0}$
return
Algorithm 7 exponential degree algorithm when $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$
C.2 Polynomial Time Algorithm
The polynomial-degree algorithm combines the three sub-algorithms in Section B.6 together to filter the observations $\mathbf{X}$ and through the weight $\omega$. The theoretical guarantee is provided in Section B.6.
1
2import $\mathbf{X},N,K,\sigma=1$
3import Algorithm 3, 4, 5, 6
4
[…]
6Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma=1,\epsilon$ as 2.3 and 2.4
7if $k_{1}^{\star}\leq k_{2}^{\star}$ then
8 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
9
10else
11 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
12
[…]
14Execute Algorithm 3, and record the returning value $\texttt{R}_{1}$
15if $\texttt{R}_{1}$ is None then
16   Reject $H_{0}$
17 return
18
19else
20 $\mathbf{Y}=\texttt{R}_{1}$
21
22Set $\omega=\mathbf{1}_{[k^{\star}]}$
23if $N>k^{\star}$ then
24   Execute Algorithm 4, and record the returning value $\texttt{R}_{2}$
25else
26   Execute Algorithm 5, and record the returning value $\texttt{R}_{2}$
27if $\texttt{R}_{2}$ is None then
28   Reject $H_{0}$
29 return
30
31else
32 $\omega=\texttt{R}_{2}$
33
34Execute Algorithm 6, and record the returning value $\texttt{R}_{3}$
[…]
37   Reject $H_{0}$
38
39else
40   Accept $H_{0}$
41
return
Algorithm 8 polynomial degree algorithm when $\rho\gtrsim P(\epsilon,k^{\star},N,1)$.

### v3

@@ 15 $\ln\left(\frac{1}{\alpha}\right)$ Dominated by the term 3
Table 2: terms from (C.23) and corresponding required conditions on $\left\lVert P^{\star}\mu\right\rVert_{2}$.
@@
$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k^{\star}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (C.24)
Once this requirement is satisfied, by Lemma C.17 and the inequality (C.23), we know that $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}P^{\star}X_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|$ will be less than $c_{1}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{0}$ and greater than $c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{1}$ for some $0<c_{1}<c_{2}<1$, and both with high probability. $c_{1}$ and $c_{2}$ are universal and their values can be determined following proper choices of $\epsilon,\alpha,N,k^{\star}$. From the argument at the end of the proof of Theorem 3.8, we know (C.24) leads to (3.6). The proof is completed. ∎
C.7 Proof of Corollary 3.15
@@
The techniques are parallel to the proof of the lower bounds above. By Lemma 4.3, for $k=k_{p,1}^{\star},c=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$, there exists a $\theta\in K$ such that $\left\lVert\theta\right\rVert_{p}=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{1}{(k_{p,1}^{\star})^{1/4}\sqrt{N}}\sigma$. Again, we impose a uniform prior on the set $\left\{\mu\left\lvert\right.\mu\in K,\mu_{i}=\kappa\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\in I\right\}$ for some positive constant $\kappa$. By the same arguments of Appendix C.2, if $\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$, or equivalently $\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}$ is sufficiently small, then there is no test with uniformly small type-I and type-II errors. We have
@@
and $\theta\in K$ from Lemma 4.3 satisfies $\left\lVert\theta\right\rVert_{p}=\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(Nk_{p,2}^{\star})^{1/4}}\sigma$. Consequently, it is not difficult to verify that $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\epsilon$. By Lemma C.1, we establish the second lower bound that $\rho_{\text{critical}}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$.
Suppose now that $\theta$ is the vector from Lemma 4.3 when $k=k_{p,3}^{\star}$. Subsequently, we know $\left\lVert\theta\right\rVert_{2}^{2}\leq\left\lVert\theta\right\rVert_{p}^{p}\cdot\left\lVert\theta\right\rVert_{\infty}^{2-p}\leq\epsilon^{2}\sigma^{2}$. Following the same reasoning in Appendix C.4, we know there always exists a strategy for $\mathcal{C}$ to discourage the testing. Therefore, we establish the third lower bound. The proof is completed. ∎
C.9 Proof of Theorem 4.6
@@
Lemma C.18 (Truncation of a $p$-convex orthosymmetric set).
@@
Lemma C.18 is straightforward from the definition. For any $\theta_{J,1},\theta_{J,2}\in K_{J}$, suppose $\theta_{J,1}=\beta_{1,J},\theta_{J,2}=\beta_{2,J},\beta_{1},\beta_{2}\in K$, then we know for any $\lambda\in[0,1]$, we have $\lambda\theta_{J,1}^{p}+(1-\lambda)\theta_{J,2}^{p}=\lambda\beta_{1,J}^{p}+(1-\lambda)\beta_{2,J}^{p}=(\lambda\beta_{1}^{p}+(1-\lambda)\beta_{2}^{p})_{J}\overset{\text{(i)}}{=}\widetilde{\beta}_{J}^{p}\in K_{J}^{p}$, where $\widetilde{\beta}\in K^{p}$ and (i) is from that $K$ is p-convex. Consequently, $K_{J}$ is p-convex.
@@
Suppose first $\epsilon\leq\frac{1}{\sqrt{N}}$. In this case, we have $\frac{j^{1/p-1/4}}{\sqrt{N}}\sigma\geq\frac{\sqrt{\epsilon}j^{1/p-1/4}}{N^{1/4}}\sigma$ for any $0\leq j\leq d$. By the monotonicity of Kolmogorov widths and the definitions of $k_{p,1}^{\star}$ and $k_{p,2}^{\star}$, we know $k_{p,1}^{\star}\leq k_{p,2}^{\star}$. Later calculations show that in this case we also have $k_{p,2}^{\star}\leq k_{p,3}^{\star}$. This implies $k_{p,1}^{\star}\leq k_{p,2}^{\star}\leq k_{p,3}^{\star}$. For $k_{p,1}^{\star}$ and its corresponding $P_{p,1}^{\star}$, without loss of generality we assume $P_{p,1}^{\star}\theta=(\theta_{1},\dots,\theta_{k_{p,1}^{\star}},0,\dots,0)$. Regard $P_{p,1}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,1}^{\star}}$, by the Hölder’s inequality B.17, we know $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ for any $\theta\in K$. Further define $k_{p,1,1}^{\star}:=\max\limits\left\{j\left\lvert\right.0\leq j\leq k_{p,1}^{\star},D_{j-1}(K)>\frac{j^{1/4}}{\sqrt{N}}\sigma\right\}$. Since $\epsilon\leq\frac{1}{\sqrt{N}}$, by the proof of Corollary 3.15 and Lemma C.18, we know that if $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, then there exists a desired test. However, this is already guaranteed since $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ and $k_{p,1}^{\star}\geq k_{p,1,1}^{\star}$.
@@
C.10 Verification of the Minimality in Section 5.2